OpenStudy (aravindg):

calculate mass of caco3 required to liberate 10 litres of co2 at stp

OpenStudy (mani_jha):

CaCO3=CaO+CO2. 10 litres of co2= 10/22.4 moles at STP. So no of moles of CaCO3=10/22.4. So, \[mass/molecular weight = 10/22.4\] Molecular weight = 100 Calculate mass

OpenStudy (aravindg):

wow thx i hav another qn too pls help